[figure: 9.7 cm]

POINT-CONTACT SPECTROSCOPY OF Ce_{1−x}RE_{x}Ni_{5} (RE = Pr, Nd)

T. Kačmárová^{a}, M. Reiffers^{b}, J. Garcia Soldevilla^{c}, J.C. Goméz Sal^{c}, J.I. Espeso^{c} and J.A. Blanco^{d} aFaculty of Sciences, University of P.J. Šafárik, Park Angelinum 9, SK-041 54 Košice, Slovakia bInstitute of Experimental Physics, Watsonova 47, SK-043 53 Košice, Slovakia

cCITYMAC, Facultad de Ciencias, Universidad de Cantabria, 39005 Santander, Spain dDepartamento de Fisica, Universidad de Oviedo, E-33007 Oviedo, Spain

We report on the study of the modification of the electron-quasiparticle interaction (EQI) function in Ce_{1−x}RE_{x}Ni_{5} (RE = Pr, Nd) pseudobinary compounds for x = 0, 0.2, 0.5, 0.8 and 1 by point-contact (PC) spectroscopy. CeNi_{5} is the Stoner enhanced paramagnet which is characteristic by the influence of spin fluctuation on its properties. On the other hand, PrNi_{5} is van Vleck paramagnet and NdNi_{5} is the ferromagnetic compound both with crystalline electric field (CEF) influence on their properties. The measured PC spectra (d^{2}V/dI^{2} - directly proportional to the EQI function) of the heterocontacts between Ce_{1−x}RE_{x}Ni_{5} and Cu in ballistic regime above and below T_{C} are presented. The intensity of spin fluctuation peak at ∼ 2.5 meV in CeNi_{5} is decreasing with increasing Nd content and it vanishes for x = 0.5. The CEF peaks from Nd ions are arising at x = 0.5 in magnetically ordered state. The main CEF peak from Pr ions appear already for x = 0.2. Its position is shifted to lower energy and it reaches finally 4 meV in PrNi_{5}. The spin fluctuation peak at ∼ 2.5 meV disappears for x = 0.5.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : T. Kačmárová

Address for correspondence : Faculty of Sciences University of P.J.Šafárik Park Angelinum 9 SK-041 54 Košice Slovakia

Email address : salonova@kosice.upjs.sk